One hundred sixty eight inches of string were used to make these squares, which are all the same size. What is the total area of the ten squares?: One hundred sixty eight inches of string were used to make these squares, which are all the same size. What is the total area of the ten squares?
Answer by nerdybill(1123)

For any square, all four sides are equal.
Let x = length of one side of the square
.
Then, the perimeter of each square is:
4x
.
If he has 10 of these squares:
10(4x) = 40x
.
Finally, since he has 168 inches of string:
40x = 168
x = 168/40
x = 4.2 inches
.
So, the area of a single square is then:
x^2
= (4.2)^2
= 17.64 sq inches
.
And, if he has 10 of these squares:
10*17.64 = 176.4 square inches